### v1

# A Simple Latent Diffusion Approach for Panoptic Segmentation and Mask Inpainting

Abstract
Panoptic and instance segmentation networks are often trained with specialized object detection modules, complex loss functions, and ad-hoc post-processing steps to handle the permutation-invariance of the instance masks. This work builds upon Stable Diffusion and proposes a latent diffusion approach for panoptic segmentation, resulting in a simple architecture which omits these complexities. Our training process consists of two steps: (1) training a shallow autoencoder to project the segmentation masks to latent space; (2) training a diffusion model to allow image-conditioned sampling in latent space. The use of a generative model unlocks the exploration of mask completion or inpainting, which has applications in interactive segmentation. The experimental validation yields promising results for both panoptic segmentation and mask inpainting. While not setting a new state-of-the-art, our model’s simplicity, generality, and mask completion capability are desirable properties. The code and models will be made available. [^1] https://github.com/segments-ai/latent-diffusion-segmentation
1 Introduction
The image segmentation task [47, 51] has gained a lot of attention in the literature, encompassing three popular subfields: semantic, instance, and panoptic segmentation. Over the years, segmentation tools have proven their usefulness for a wide range of applications, such as autonomous driving [19], medical imaging [50], agriculture [17], and augmented reality [1, 26]. Current methods are built upon convolutional networks [28] and transformers [73, 21] to learn hierarchical image representations, and to simultaneously leverage large-scale datasets [85, 45]. Earlier segmentation approaches relied on specialized architectures [29, 7, 67, 75], such as region proposal networks [58] and dynamic convolutions [34]. More recent approaches [16, 8] advocate an end-to-end strategy but introduce complex loss functions, i.e., bipartite matching using object queries to separate instances. Some works have shown promising results without the necessity for labels [70, 71, 72, 18, 76, 77], but likewise require highly-specialized modules, such as region proposal networks or clustering. Differently, we seek to leverage generative models to bypass the aforementioned components.
This observation aligns with recent works [49, 12, 41, 10, 11] advocating for general computer vision models in and attempt to unify the field. These pioneering works limit the adoption of task-specific components and instead use a generative process, such as autoregressive modeling, rendering its design applicable to different tasks and/or modalities. In a similar vein, we take inspiration from recent text-to-image diffusion models [59, 57, 52, 20, 31, 32, 56] to tackle the segmentation task in a generative fashion. In addition to the diffusion model’s general architecture, we further motivate this decision as follows: (i) diffusion models are capable of generating images with high photorealism and diversity; (ii) they perform on par with autoregressive priors while being more computationally efficient [63, 57]; (iii) they learn high-quality spatial representations, advantageous for dense prediction tasks; (iv) they naturally exhibit image editing capabilities. We build further upon these properties by harnessing the generative power of Latent Diffusion Models (LDMs) [59] for image segmentation applications.
Figure 1: We present a simple generative approach for the segmentation task that builds upon Stable Diffusion [59]. The key idea is to leverage the diffusion process to bypass complex detection and segmentation modules, and to unlock mask inpainting. The generative process is conditioned on RGB images to predict panoptic masks in an iterative fashion.
To realize this objective, we introduce LDMSeg, a simple Latent Diffusion Model for Segmentation, visualized in figure 1. In summary, our contributions are threefold:
  • Generative Framework: This paper proposes a fully generative approach based on Latent Diffusion Models (LDMs) for panoptic segmentation. Our approach builds upon Stable Diffusion [59] to strive for simplicity and to ease compute. We first study the class-agnostic setup to liberate panoptic segmentation from predefined classes [40].
  • General-Purpose Design: Our approach circumvents specialized architectures, complex loss functions, and object detection modules, present in the majority of prevailing methods. Here, the denoising objective omits the necessity for object queries, region proposals, and Hungarian matching [42]. This simple and general approach paves the way for future extensions to a wide range of dense prediction tasks, e.g., depth prediction, saliency estimation, etc.
  • Mask Inpainting: We successfully apply our approach to scene-centric datasets and demonstrate its mask inpainting capabilities for different sparsity levels.
2 Related Work
Panoptic Image Segmentation.
Panoptic segmentation [38] has lately gained popularity as it combines semantic and instance segmentation. In particular, its goal is to detect and segment both stuff-like (e.g., vegetation, sky, mountains, etc.) and thing-like (e.g., person, cat, car, etc.) categories. Earlier works modified instance segmentation architectures to additionally handle (amorphous) stuff categories as they are hard to capture with bounding boxes. For instance, Kirillov et al. showed promising results by making independent predictions using semantic and instance segmentation architectures [38], and later by integrating a semantic segmentation branch with a feature pyramid network (FPN) into Mask R-CNN [37]. Other works [79, 14] extend this idea by relying on specialized architectures and loss functions from both segmentation fields. More recent works [8, 74, 16, 15, 81, 84] handle things and stuff categories in a unified way via object queries and Hungarian matching [42]. These works generally depend on specialized modules – such as anchor boxes [58], non-maximum suppression [33, 6], merging heuristics [38, 79], or bipartite matching algorithms [8, 16, 74], etc. – to generate panoptic masks. Instead, we propose a task-agnostic generative framework to bypass these components. Consequently, we refrain from using task-specific augmentations, such as large-scale jittering or copy-paste augmentations [25].
General-purpose Frameworks.
Similar to our work, a number of task-agnostic solutions have been suggested that cast vision problems as a generative process. Kolesnikov et al. [41] minimized task-specific knowledge by learning task-specific guiding codes, using a language model, to train a separate vision model. Chen et al. [11] simplified this procedure by framing several vision tasks as language modeling tasks, within a single model. Lu et al. [49] also follow this route and show promising results for a large variety of vision and language tasks using a unified framework. Each work presents the input as a sequence of discrete tokens which are subsequently reconstructed via autoregressive modeling. Other works [78, 4] leverage masked image modeling to train a single model for multiple vision tasks. Differently, we leverage the denoising process in continuous latent space, which is well-suited to handle dense prediction tasks with high-dimensional inputs [32, 12, 59].
Denoising Diffusion Models.
Denoising diffusion models [62, 32, 63] were introduced as a new class of generative models. Recent strategies [20, 52, 31, 57] additionally leverage text as guidance, e.g., via CLIP embeddings, to achieve results with impressive realism and control. Building upon its success, a few diffusion-based solutions appeared in the segmentation literature. However, they have undesirable properties: (i) the inability to differentiate between instances in an image [2, 3, 5], (ii) the necessity for specialized architectures and loss functions [27, 80], or (iii) the dependence on object detection weights and bit diffusion [12, 13]. The closest related work from Chen et al. [12], presents a framework for panoptic segmentation by leveraging the diffusion process in pixel-space. In contrast, we rigorously follow latent diffusion models by relying on continuous latent codes without the necessity for object detection modules. Importantly, we can effortlessly leverage public image-diffusion weights [59] as we keep the architecture task-agnostic.
3 Method
Figure 2: Overview of LDMSeg. Inspired by latent diffusion models, we present a simple diffusion framework for segmentation and mask inpainting. The approach consists of two stages: (i) learn continuous codes $z_{t}$ with a shallow autoencoder on the labels (Sec. 3.1); (ii) learn a denoising function conditioned on image latents $z_{i}$ (Sec. 3.2). In the second stage, the error between the predicted noise $\hat{\epsilon}$ and the applied Gaussian noise $\epsilon$ is minimized. During inference, we traverse the denoising process by starting from Gaussian noise. The models $f_{t}$ and $f_{i}$ respectively encode the labels and images. While we rely on the image encoder $f_{i}$ from Stable Diffusion [59], we focus on $f_{t}$ and $g$ for segmentation. Our aim is to prioritize generality by limiting task-specific components.
Preliminaries.
The aim of this paper is to train a fully generative model for panoptic segmentation via the denoising diffusion paradigm [32]. While this generative approach results in longer sampling times than discriminative methods, we justify this decision through four different lenses: (i) ease of use – diffusion models omit specialized modules, such as non-maximum suppression or region proposal networks, are stable to train, and exhibit faster sampling than autoregressive models [68]; (ii) compositionality – this generative approach captures complex scene compositions with high realism and diversity, while also enabling image editing [57]; (iii) dataset-agnostic – we rely on spatially structured representations that are not tied to predefined classes or taxonomies; (iv) computational cost – latent diffusion models [59] reduce the computational requirements by modeling latents instead of pixels with an autoencoder [36, 69]. Motivated by these points, we create a framework that builds upon latent diffusion models’ generative power [59, 57] for the segmentation task.
Problem Setup.
The considered segmentation task requires a dataset of images $\mathcal{X}=\{x_{1},\ldots,x_{n}\}$ and corresponding ground truth panoptic segmentation masks $\mathcal{Y}=\{y_{1},\ldots,y_{n}\}$. We don’t make any distinction between things or stuff classes but handle all classes identically. Particularly, we first only predict the individual panoptic IDs and not the individual classes, rendering our approach class-agnostic. We will also show results in the experiments when predicting class-labels.
Assume that all images are of size $H\times W$. We train a segmentation model to realize the mapping $\mathbb{R}^{3\times H\times W}\rightarrow\mathbb{R}^{N\times H\times W}$. For each pixel the model performs a soft assignment over the instances $\{1,\ldots,N\}$. Let the latent representations $z_{i}$ and $z_{t}$ respectively refer to the image and target features after the projection to $D$-dimensional latents $\mathbb{R}^{D\times H/f\times W/f}$, where $f\in\mathbb{N}$ denotes the resizing factor.
From a high-level perspective, our method has two key components: learning the prior over segmentation latents $z_{t}$ and learning a conditional diffusion process in latent space. First, we train a shallow autoencoder to capture the prior distribution $p(z_{t})$ that learns to compress the labels into compact latent codes $z_{t}$. Second, we train a diffusion process – conditioned on the image and target features $p(y|z_{t},z_{i})$. This component is responsible for decoding noisy target features, guided by image features $z_{i}$. We make a similar derivation as [57] but condition the generative process on images. Formally, this two-step procedure allows us to construct the conditional distribution $p(y|x)$ via the chain rule:
$p(y|x)=p(y|z_{t},z_{i})\cdot p(z_{t}).$ (1)
Both terms $p(z_{t})$ and $p(y|z_{t},z_{i})$ are reflected in our framework as two separate training stages (see figure 2). Section 3.1 discusses the learning of the prior via autoencoding and section 3.2 leverages this prior to train a latent diffusion model for segmentation.
3.1 Stage 1: Compress Segmentation Targets
In the first step, we train a network to compress the task-specific targets into latent codes. While we will focus on panoptic segmentation masks, we note that a similar strategy is applicable to other dense prediction targets, such as depth or surface normals estimation.
Motivation.
The motivation to design a shallow autoencoder stems from the observation that segmentation maps differ fundamentally from images as they are lower in entropy. First, these masks typically contain only a small number of unique values as they only capture the object’s general shape and location in the scene. Second, neighbouring pixels are strongly correlated and often identical, resulting in largely spatially redundant information. We conclude that the segmentation task only necessitates a shallow network to efficiently compress the task’s targets and to reliably capture the prior distribution $p(z_{t})$. Interestingly, this conclusion holds for a myriad of dense prediction tasks, such as panoptic segmentation, depth prediction, saliency estimation etc. This also justifies why we refrain from using more advanced autoencoders that rely on computationally demanding architectures with adversarial or perceptual losses [35, 83], e.g., VQGAN [23], typically used to encode images [59].
Encoding.
We analyze several encoding strategies to represent the input segmentation map $y$, latents $z_{t}$ and output $\hat{y}$.
1. Input $y$: Let $N$ denote the maximum number of instances per image. RGB-encoding (3 channels), bit-encoding ($\log_{2}N$ channels), one-hot-encoding ($N$ channels), or positional-encoding [73] are all justifiable options. We follow [12] in representing the instance IDs as bits, which is a natural choice given discrete segmentation maps.
2. Latent code $z_{t}$: No vector quantization or dimensionality reduction is applied to the latent codes. We empirically found that directly using the continuous latents resulted in a simple architectural design with a small memory footprint (shallow) while simultaneously demonstrating encouraging segmentation results for both reconstruction and generation.
3. Output $\hat{y}$: The output is one-hot encoded to train the autoencoder with a standard cross-entropy loss. In practice, we define $N$ output channels to reconstruct $N$ different instances in a scene, e.g., $128$.
Architectural Design.
The architecture of the autoencoder follows a simple and shallow design. It comprises an encoder $f_{t}$ and decoder $g$, yielding the overall function $g\circ f_{t}$. The encoder $f_{t}$ includes only a few strided convolutions to compress the targets and is inspired by ControlNet [82]. For instance, targets with size 512$\times$512 can be resized efficiently to 64$\times$64, in order to leverage the latent space of Stable Diffusion [59] by stacking 3 convolutions of stride 2. Similarly, the decoder $g$ consists of one or more transpose convolution to upscale the masks and minimize a loss at pixel-level. As a consequence, the number of trainable parameters is at least 2 orders of magnitudes smaller than the amount of parameters in the diffusion model $h$ ($\approx 2$M vs. $800$M). The model’s shallow design brings several advantages to the table: fast training, good generalization across datasets, and applicable to inpainting without architectural changes.
Loss Function.
The autoencoder aims to minimize the reconstruction error $\mathcal{L}_{\text{rec}}$ between the outputs $\hat{y}$ and the one-hot encoded segmentation masks $y$. While regression losses are a valid choice, we opt for a categorical loss due to segmentation task’s discrete nature. The reconstruction loss comprises two loss terms: (i) the cross-entropy loss $\mathcal{L}_{\text{ce}}$ enforces unique and confident predictions for each pixel; (ii) the mask loss $\mathcal{L}_{\text{m}}$ further refines the segmentation masks by treating each instance individually, alleviating the need for exhaustively labeled images. This term is implemented via the BCE and Dice losses [66, 29]. Note that this autoencoding strategy prevents the need for Hungarian matching [42].
Typically, latent diffusion models incorporate a penalty term $\Omega$ into the loss formulation to align the bottleneck latents with a standard Gaussian distribution $\mathcal{N}(0,I)$. This term generally takes the form of a KL divergence, resulting in a variational autoencoder [36]. However, we empirically observed that weight decay regularization suffices to keep the weights $w$, and by extension the latents $z_{t}$, bounded. Consequently, the magnitude of the model’s weights $||w||_{2}$ is penalized, resulting in the final loss:
$\begin{split}\mathcal{L}_{\text{AE}}(w;y)&=\mathcal{L}_{\text{rec}}(w;\hat{y},y)+\Omega(z_{t},w)\\ &=\mathcal{L}_{\text{ce}}(w;\hat{y},y)+\mathcal{L}_{\text{m}}(w;\hat{y},y)+\lambda||w||_{2}^{2},\end{split}$ (2)
where $\hat{y}$ denotes the reconstructed segmentation map. Furthermore, we follow PointRend [39] to select logits that correspond with uncertain regions. This strategy limits the memory consumption as well as the total training time.
3.2 Stage 2: Train a Denoising Diffusion Model
Image-Conditioned Diffusion Process.
Figure 3: Diffusion Process and SNR. (1) During training we randomly sample a timestep from $[1,T]$ in the denoising process. We can increase the RGB image’s importance by strengthening the noise: (i) Following [59, 12], we downscale the latents $z_{c}$ using scaling factor $s\in\mathbb{R}$ and demonstrate its impact – Row 1 ($s=1.0$) is clearly easier to decode than row 2 ($s\approx 0.18$ [59]). (ii) Losses for timesteps near 00 are further downscaled to avoid overfitting. Both strategies enforces the model to focus on the RGB image in generating plausible segmentation maps. Note that we don’t apply explicit constraints to the prior distribution $p(z_{t})$, e.g., match a standard Gaussian $\mathcal{N}(0,1)$. (2) During sampling the denoising process is traversed from right to left in $T_{s}$ iterations.
The second stage of the framework models the function $h$ via a conditional diffusion process. In particular, we aim to learn a diffusion process over discrete timesteps by conditioning the model on images and noisy segmentation masks for each individual timestep. We further follow the formulation of Stable Diffusion [59]. Hence, this process is carried out in a joint latent space – using our trained segmentation encoder $f_{t}$ and given image encoder $f_{i}$ [59] – in order to project the targets and images to their respective latent representations $z_{t}$ and $z_{i}$. We discuss the training and inference procedures next.
During training we linearly combine the noise $\epsilon$ with the latents $z_{t}$ for a randomly sampled timestep $j\in[1,T]$:
$\tilde{z}^{j}_{t}=\sqrt{\bar{\alpha}_{j}}z_{t}+\sqrt{1-\bar{\alpha}_{j}}\epsilon,$ (3)
where $\bar{\alpha}_{j}$ is defined by the noise schedule, following Rombach et al. [59]. The latents are subsequently fused $\{\tilde{z_{t}}^{j},z_{i}\}$ via channel-wise concatenation as $z_{c}\in\mathbb{R}^{2D\times H/f\times W/f}$, before feeding them to the UNet [60, 59] $h_{\theta}$, parameterized with weights $\theta$. Despite the examination of various fusing techniques (i.e., fusing intermediate features via cross-attention or via modality-specific branches) we found that straightforward concatenation at the input works surprisingly well. At the output, the reconstruction error between the added Gaussian noise $\epsilon\sim\mathcal{N}(0,I)$ and the predicted noise $\hat{\epsilon}$ is minimized [32] as
$\begin{split}\mathcal{L}_{\text{LDMSeg}}(\theta;\epsilon)&=\mathbb{E}_{z_{c},\epsilon\sim\mathcal{N}(0,I),j}\left[||\epsilon-h_{\theta}(z_{c},j)||_{2}^{2}\right],\end{split}$ (4)
where each timestep $j$ is uniformly sampled from $[1,T]$. To reduce training time, we downscale the loss for small timesteps, i.e. $j<25\%\cdot T$. These latents have a high signal-to-noise ratio (SNR) and are thus relatively easy to denoise without modeling semantics. We finally refer to algorithm 1 for an overview of the forward pass during training.
Algorithm 1 Forward pass in Pytorch.
# f_i, f_t: image encoder, segmentation encoder
# h: denoising UNet
# x: images of size [bs, 3, H, W]
# y: bit maps of size [bs, log(N), H, W]
# s, T: scaling factor, number of train steps
y = 2 * y - 1      # bit maps to range [-1, 1]
x = x / 127.5 - 1   # images to range [-1, 1]
z_t = f_t(y) * s    # encode bit maps and rescale
z_i = f_i(x) * s    # encode images and rescale
j = torch.randint((bs,) 0, T)
noise = torch.randn_like(z_t)
z_t = scheduler.add_noise(z_t, noise, j)  # Eq. 3
z_c = torch.cat([z_t, z_i], dim=1)
noise_pred = h(z_c, j)
loss = torch.sum((noise_pred - noise)**2, dim=[1,2,3])
loss = torch.mean(loss * scheduler.weights[j])
 scheduler.weights: attribute - array of length $T$ with loss weights
Algorithm 1 Forward pass in Pytorch.
Algorithm 2 Sampling process in Pytorch.
# f_i, g: image encoder, segmentation decoder
# h: denoising UNet
# x: image of size [3, H, W]
# T, Ts: number of training steps, inference steps
x = x / 127.5 - 1  # image to range [-1, 1]
z_i = f_i(x) * s    # encode image and rescale
z_t = torch.randn_like(z_i)  # Gaussian noise
for j in scheduler.inference_steps:
    z_c = torch.cat([z_t, z_i], dim=0)
    noise_pred = h(z_c, j)
    j_prev = j - T // Ts
    # apply Eq. 5
    z_t = scheduler.step(z_t, noise_pred, j, j_prev)
y_pred = g(z_t)  # decode segmentation latents
  scheduler.inference_steps: attribute - chosen sampling timesteps
scheduler.step: function - predict previous sample
Algorithm 2 Sampling process in Pytorch.
During sampling, the denoising process is traversed from right to left (see figure 3). It starts from Gaussian noise and progressively adds more details to the segmentation map as controlled by the input image. We rely on the DDIM scheduler [63] to apply this denoising process over a small number of sampling timesteps $T_{s}<<T$. Here, the previous sample is computed as
$\tilde{z}_{t}^{j-1}=\frac{\sqrt{\bar{\alpha}_{j-1}}}{\sqrt{\bar{{\alpha}}_{j}}}(\tilde{z}_{t}^{j}-\sqrt{1-\bar{\alpha}_{j}})\cdot\hat{\epsilon}+\sqrt{1-\bar{\alpha}_{j-1}}\cdot\hat{\epsilon},$ (5)
which follows directly from equation 3. Recall that this strategy allows us to model $p(y|z_{t},z_{i})$, as defined in equation 1. Finally, algorithm 2 provides the details of the sampling process and figure 3 serves as an illustration.
Image Descriptors.
Complementary, we briefly experimented with adding guidance via Stable Diffusion’s cross-attention layers to improve sample quality. This mechanism was initially intended for text embeddings, but can be used to feed any image descriptor. We didn’t observe improvements with self/weakly-supervised priors, like image embeddings from CLIP [55], or when using captions from BLIP [43]. We currently bypass the cross-attention layers with skip connections, but this is a promising direction to explore further.
3.3 Segmentation Mask Inpainting
Setup.
Our image-conditioned diffusion model is well-suited to complete partial segmentation masks. This is potentially useful for (i) completing sparse 2D segmentation masks obtained from projecting 3D point cloud segmentation labels onto images; (ii) refining noisy segmentation labels that were propagated from one frame to the next using optical flow; (iii) interactive image labeling with rough brush strokes. In contrast to image editing [59], the considered segmentation maps contain empty regions, which we initialize with 00’s. We simulate different sparsities in section 4.2.
Existing state-of-the-art approaches are not designed for these applications as they decouple the output from the input via bipartite matching [16, 8]. They require additional steps to match predictions with the given partial segmentation IDs (e.g., via majority voting). In contrast, diffusion models act on the corrupted masks by default.
Inpainting Process.
Ideally, we can tackle inpainting problems out-of-the-box, i.e., without finetuning. To achieve this goal, the inference loop, previously discussed in algorithm 2, is modified. Assume that we have a dataset of pairs $(y,m)$. Each pair contains a sparse segmentation mask $y\in\{0,1\}^{\log_{2}N\times H\times W}$, represented as a bitmap, and a valid mask $m\in\{0,1\}^{H\times W}$, represented as a boolean mask. Now, the diffusion process should fill in the missing regions in $y$, determined by the zeros in $m$. At each step of the denoising process, the latents corresponding to the given pixels in $m$ are fixed. The key differences with the sampling process are highlighted in algorithm 3.
Algorithm 3 Inpainting process in Pytorch.
# f_i, f_t: image encoder, segmentation encoder
# g, h: segmentation decoder, denoising UNet
# x: image of size [3, H, W]
# y: sparse bit map of size [log(N), H, W]
# m: boolean mask indicating valid pixels, size [H, W]
# T, Ts: number of train steps, inference steps
y = 2 * y - 1     # bit maps to range [-1, 1]
 y[:, $\sim$ m] = 0  # set invalid regions to 0
x = x / 127.5 - 1  # image to range [-1, 1]
z_i = f_i(x) * s    # encode image and rescale
 z_t_masked = f_t(y) * s  # encode bit map and rescale
 m = interpolate(m, size=z_i.shape[-2:])
z_t = torch.randn_like(z_i)
for j in scheduler.inference_steps:
    z_c = torch.cat([z_t, z_i], dim=0)
    noise_pred = h(z_c, j)
    # apply Eq. 3 (solve for z_t)
     z_t = scheduler.remove_noise(z_t, noise_pred, j)
     z_t[:, m] = z_t_masked[:, m]  # keep latents
    j_prev = j - T // Ts
    # apply Eq. 3
     z_t = scheduler.add_noise(z_t, noise_pred, j_prev)
y_pred = g(z_t)  # decode segmentation latents
  scheduler.inference_steps: attribute - chosen sampling timesteps
scheduler.remove_noise: function - remove noise from sample
scheduler.step: function - predict previous sample
Algorithm 3 Inpainting process in Pytorch.
4 Experiments
Figure 4: Qualitative Results on COCO. The figure displays results on COCO val2017. We follow the inference setup (section 3.2) to sample from our model. Only the $\operatorname*{arg\,max}$ operator is applied for post-processing. Key observation: our model disentangles overlapping instances in challenging scenes without complex modules or post-processing. To visualize, segments are assigned to random colors, and missing (VOID) pixels in the ground truth are black.
Figure 5: Qualitative Results on Cityscapes. The figure displays predictions on Cityscapes’s validation set. Key observation: the model produces panoptic masks covering various visual concepts. Thin structures, e.g., poles, are difficult to capture due to the diffusion process in latent space.
Figure 6: Mask Inpainting. The figure visualizes generated samples for different granularity levels by following section 3.3. Key observation: the model is able to fill in missing regions by propagating the partially given segmentation IDs using an image-conditioned diffusion process. Global mask inpainting (left) works out-of-the-box while sparse mask inpainting (right) shows inaccuracies. We hypothesize that this can be addressed by further finetuning LDMseg on sparse inpainting masks.
Figure 7: Ablation Sampling Steps. The figure shows the impact of the number of sampling steps on the PQ, SQ, and RQ metrics (COCO val).
Figure 8: Ablation KL Loss. The figure shows the impact of a KL loss on the PQ, SQ, and RQ metrics (COCO val).
Figure 7: Ablation Sampling Steps. The figure shows the impact of the number of sampling steps on the PQ, SQ, and RQ metrics (COCO val).
Figure 8: Ablation KL Loss. The figure shows the impact of a KL loss on the PQ, SQ, and RQ metrics (COCO val).
Figure 9: Sparsity Experiment. The figure shows the PQ for different drop rates and block sizes ($B\times B$ pixls) in mask $m$ (COCO val).
Figure 10: Impact of Gaussian Noise. We show predictions on COCO val2017 when sampling different noise maps. The images and the generated masks are visualized in row 1 and rows 2 & 3 respectively. All predictions in the same row were generated from the same noise, visualized in column 1. Key observation: different initializations generate panoptic masks with different IDs (see row 2 vs. row 3). The Gaussian noise implicitly functions as object queries [15, 8] or region proposals [29].
4.1 Experimental Setup
Dataset.
We conduct the bulk of our experiments on COCO [45]. The standard train and val sets are adopted for training (118k images) and evaluation (5k images) respectively. During training, we use the panoptic masks – containing both stuff and things classes. For the class-agnostic setup, we only aim to predict instances, without their categories.
Architecture and Training Setup.
The maximum number of detectable segments $N$ is set to $128$. We resize the input to $512\times 512$, apply random horizontal flipping, and randomly assign segmentation IDs to integers in [$10,N-$ 1]. The segmentation encoder $f_{t}$ processes the panoptic mask $y$ by leveraging 3 convolutional layers with stride 2 and SiLU [22] activations. This results in a resizing factor $f$ of $8$ and latents with size $4\times 64\times 64$. We rely on the image encoder $f_{i}$ from Rombach et al. [59] (VAE) to convert the image $x$ to latents of the same size. We adopt Stable Diffusion’s pretrained weights to initialize the UNet $h$, and its rescaling factor $s$ to lower the SNR as $s\cdot z_{c}$ [59]. Notably, $4$ zero-initialized channels are appended to the first convolutional layer of $h$, allowing us to operate on the concatenated input $z_{c}$. We additionally incorporate self-conditioning [13] to improve sample quality. Our segmentation decoder $g$ consists of 2 transpose convolutions, resulting in an upscaling factor of $4$. We randomly sample $j$ from $1000$ discrete timesteps and linearly decay the loss for the bottom $25$% to lower the impact of samples with high SNR. The AdamW [48] optimizer is adopted with a learning rate of $1e^{-4}$ and weight decay of $5e^{-2}$. Finally, the first stage is trained for 60k iterations with a batch size of $16$ while the second stage is trained for 50k iterations with a batch size of 256 unless stated otherwise.
Inference Setup and Evaluation Protocol.
During inference, the DDIM scheduler [64] generates samples with $50$ equidistant timesteps in latent space. At the end of the denoising process, we decode and upscale the segmentation logits (output of $g$) with a factor of $2$ using bilinear interpolation. The $\operatorname*{arg\,max}$ operator produces the final (discrete) per-pixel segmentation masks. We benchmark our approach with the Panoptic Quality (PQ) evaluation protocol, defined by Kirillov et al. [38]. PQ is the product of two quality metrics: the segmentation quality which measures the intersection-over-union of matched segments (IoU) and the recognition quality which measures the precision and recall (F-score).
4.2 Panoptic Segmentation Results
Image-conditioned Mask Generation.
We visualize predictions of LDMSeg in figure 4 and 5 on a few validation images from COCO [45] and Citysapes [19] respectively. The model divides an image into semantically meaningful groups and captures different objects in the scene, e.g., persons, horses, cats, cars, etc. Notably, LDMSeg directly generates non-overlapping instance masks, required for panoptic segmentation. Additionally, figure 10 visualizes the predictions when sampling from our generative model (see first column) with different seeds. Different Gaussian noise maps lead to different segmentation IDs. Compared to other frameworks, the sampled noise resembles the object queries in Mask2Former and DETR [16, 8], or the regions proposals in Mask R-CNN [29].
Mask Inpainting.
Figure 6 demonstrates the inpainting performance for two different granularity levels in mask $m$. We mask random regions in the ground truth to simulate global and local completion tasks. The results demonstrate that LDMSeg is able to fill panoptic masks via the diffusion process, and without requiring additional components.
Further, figure 9 displays the PQ metric for a wide range of drop rates $d$ and block sizes $B$. While our model performs well for global inpainting, there is still room for improvement when considering very sparse inpainting tasks (e.g., dropping $16\times 16$ pixels with a drop rate of $90\%$). Particularly, for drop rates above $70\%$ and small block sizes, the reduction rate (negative slope) in PQ increases. Further improvements can be obtained by finetuning or modifying the autoencoder. For instance, shallower encoders $f_{t}$ could adapt better to sparse inputs, in exchange for lower reconstruction quality.
  Method Backbone PQ $\uparrow$
Specialist approaches:
  MaskFormer [15] ViT [21] 54.1
  Mask2Former [16] ViT [21] 56.5
  ${}^{\star}$ [16] ViT [21] 59.0
Generalist approaches:
  ${}^{\dagger}$ (Ours) UNet [60] 50.9
   LDMSeg (Ours) UNet [60] 50.8
Table 1: Comparison with Specialists (class-agnostic). The results are reported on COCO val2017. We differentiate between specialist and generalist methods and train for 50k iterations. Training details for Mask2Former are provided in the supplementary since we use a simplified version, similar to SAM [40] with MAE [30] or DINOv2 [53] initialization (see normal-$\star$). (normal-$\dagger$) denotes the adoption of Stable Diffusion’s VAE with 84M trainable parameters, resulting in 20×20$\times$20 × more parameters compared to our standard setup (84M vs. 2M) for the same performance (see section 3.1).
  Method Backbone PQ $\uparrow$
Specialist approaches:
  PanopticFPN [37] ResNet [28] 44.1
  DETR [8] ResNet [28] 45.6
  MaskFormer [15] ResNet [28] 46.5
  Mask2Former [16] Swin-L [46] 57.8
Generalist approaches:
  Painter [78] ViT [21] 41.3
  UViM [41] ViT [21] 45.8
  Pix2Seq-$\mathcal{D}^{\dagger}$ [12] ResNet [28] 50.3
   LDMSeg (Ours) UNet [60] 43.3
Table 2: State-of-the-art Comparison. The table presents the panoptic segmentation results on COCO val2017. (LDMSeg vs. specialists) Our approach almost matches the performance of several specialists on the panoptic segmentation task. (LDMSeg vs. generalists) LDMSeg outperforms Painter, but not Pix2Seq-$\mathcal{D}$ and UViM. Note that Pix2Seq-$\mathcal{D}$ leverages additional annotated training data (normal-$\dagger$) from Objects365 [61] and UViM is sensitive to the adopted guiding codes, necessitating special mechanisms like code dropout and optimal code length [41]. In contrast, we can perform mask inpainting out-of-the-box.
State-of-the-art Comparison.
Table 1 compares LDMSeg with prior art for the class-agnostic setup. Our model obtains 50.8% PQ on COCO’s validation set. However, there is still a gap with specialized methods, i.e., MaskFormer [15, 16] which obtains 59.0% PQ.
To compare with state-of-the-art methods, we train our method for 200k iterations, as well as a classifier with a simple segmentation head [9] on top of the proposals. We compare the obtained panoptic segmentation in table 2. Our generative approach can almost match several specialized approaches ($43.3\%$ vs. $44.1\%$[79, 37]. LDMSeg outperforms Painter [78] while not requiring different encoding strategies for things and stuff categories or NMS ($43.3\%$ vs. $41.3\%$). While UViM [41] obtains better results, it’s arguably less robust due to its strong reliance on specific code lengths and code dropout, in order to train its autoregressive language model (LM). In contrast, we leverage latent codes centered around zero using a shallow autoencoder, which benefits robustness. To increase the denoising difficulty, we simply drop the scaling factor $s$ (see section 3.2). Unsurprisingly, Pix2Seq-$\mathcal{D}$ [12] achieves higher PQ by adopting additional training data [61] and a pixel-level objective. We expect that larger datasets and a higher resolution will help in closing this gap.
In contrast to prior art [41, 78], LDMSeg can tackle the task of mask inpainting out-of-the-box, and without specific processing techniques. As UViM [41] is not capable of mask inpainting by design, it’s unclear how Painter [78] and Pix2Seq-$\mathcal{D}$ [12] should be modified: (i) Painter [78] relies on a predefined color palette that differs for things and stuff classes. Due to this fixed encoding (e.g., center of mass for things categories), it is unable to complete the given (random) segmentation IDs. Furthermore, its hefty post-processing is not guaranteed to preserve the IDs in the partial input mask, which is detrimental for mask inpainting. Our method is conditioned on latents that are not confined to a task-specific encoding, which preserves the given input masks. (ii) Future research could explore if our inpainting strategy (see section 3.3) can be extended to Bit Diffusion [13], used in Pix2Seq-$\mathcal{D}$ [12]. However, the unreleased weights poses a challenge in adopting this paradigm.
Ablation Studies.
We ablate the number of inference timesteps in figure 9 and observe that the PQ metric starts plateauing at 20 iterations. Longer inference schedules can further improve the recognition quality (i.e., a reduction in false positives and negatives), while keeping the segmentation quality constant ($\sim 80\%$). We report our best results with 50 iterations.
Figure 9 demonstrates the impact of adding a KL loss to the autoencoding objective in equation 2. It aligns the latents with a standard Gaussian distribution. Unsurprisingly, increasing its loss weight beyond $1e^{-5}$ is not beneficial as it hurts the reconstruction quality. We conclude that weight decay suffices to keep the segmentation latents bounded.
Table 1 verifies the hypothesis that panoptic masks don’t require a powerful VAE. In particular, we finetuned Stable Diffusion’s VAE [59] instead of our shallow autoencoder, as discussed in section 3.1. However, this results in the same final PQ (50.9% vs. 50.8%). Our shallow autoencoder contains 20×20$\times$20 × less parameters, which reduces the training time of stage 2 by 20% (see section 3.2). This observations likely holds for other vision tasks as well.
5 Conclusion
In summary, this paper presented LDMSeg, a simple yet powerful latent diffusion approach for segmentation and inpainting. It builds upon Stable Diffusion [59] to learn an image-conditioned diffusion process for segmentation in latent space. The experimental results are promising: (i) our framework bypasses specialized object detection modules, such as region proposal networks and bipartite matching, to generate panoptic masks; (ii) our model unlocks sparse panoptic mask completion without finetuning. Hence, we hope that this work will spark further interest into designing general-purpose approaches for dense prediction tasks. Due to the simple and general design of LDMSeg, there is still room for improvement in terms of accuracy and sampling speed. Evident future directions include: training the presented method on larger datasets and extending it to other dense prediction tasks, e.g., depth prediction.
A Supplementary Material
We discuss the implementation details in section B, additional results in section C, limitations in section D, and broader impact in section E.
B Implementation Details
Model Card.
Our best model is trained for 200k iterations on COCO with mixed precision training on 8 $\times$ 40GB NVIDIA A100 GPUs using the Google Cloud. We rely on the pretrained Stable Diffusion [59] weights provided by Hugging Face [24]. We also adopt its settings for the noise scheduler. The code is developed in Pytorch [54] and will be made available as well as our models.
Baseline.
Mask2Former [16] is a specialized segmentation framework that produces excellent results for the panoptic segmentation task. We follow the training recipe from ViTDet [44] and SAM [40] to leverage plain ViT backbones [21] with MAE pretrained weights [30]. Specifically, the model consists of the vision transformer backbone, a shallow neck, and a mask decoder. The latter contains $6$ masked attention decoder layers and $128$ object queries, following [16]. The loss requires Hungarian matching [42] to handle the permutation invariance of the predictions during training. To report the results, we follow its post-processing strategy to combine the classification and mask branches. We adopt the same augmentations as in the main paper, i.e., square resizing and random horizontal flipping. This baseline strikes a good balance between performance, complexity, and training speed. Additionally, we provide results by relying on the backbone and pretrained weights of DINOv2 [53], as we found this to outperform MAE pretrained weights for a ViT-B backbone. We train these models with a batch size of $32$ and a learning rate of $1.5e^{-4}$ for $50$k iterations on 8 $\times$ 16GB V100 GPUs.
Evaluation Procedure.
Our model produces excellent predictions by simply relying on the $\operatorname*{arg\,max}$ operator. No additional processing is used for the visualizations (see row 3 in figures S2 and S3). In order to report the final PQ metric, however, we eliminate noise by thresholding the predictions at $0.5$ (after applying softmax) and filtering out segments with an area smaller than $512$. These results are shown in the last row of figures S2 and S3. Notice that Mask2Former’s training objective does not impose exclusive pixel assignments, hence it needs additional post-processing steps.
C Additional Results
We show the panoptic segmentation results on COCO val2017 [45] in figure S1 using 50 timesteps. Additionally, we show (class-agnostic) masks in figures S2 and S3. The input images are resized to $3\times 512\times 512$ during training and the diffusion process acts on latents of size $4\times 64\times 64$. To visualize the masks, we assign each segment to a random color. Overall, we conclude that the model is capable of generating high-quality panoptic masks.
Figure S1: Panoptic Segmentation - Qualitative Results. The figure displays the panoptic segmentation for several images in the COCO val set.
Figure S2: Examples on COCO (1). The figure displays the generated masks on the COCO val set.
Figure S3: Examples on COCO (2). The figure displays more generated masks on the COCO val set.
Figure S4: Results for different timesteps. The figure visualizes the image-conditioned samples for the timesteps 1, 5, 10, 20, and 50 in the diffusion process. Longer sampling is required to capture more details, which is beneficial for complex scenes (e.g., cars in the background in column 4).
Figure S4 displays the results for different timesteps in the denoising process. Longer sampling benefits the generation of details, such as capturing small objects in the background or an object’s edges. This approach necessitates 10 to 50 iterations to produce high-quality segmentation masks, which is in line with latent diffusion models for images [59]. Furthermore, as the model was forced to distinguish between different instances during training, it’s unlikely that different instances will be grouped during inference. Interestingly, the model iteratively improves the predictions while not reinforcing mistakes during the generative process.
Table 1 provides the inference times for different sampling durations. Unsurprisingly, diffusion models are at least an order of magnitude slower than specialized methods (compare 0.47s vs. 0.02s). However, recent research [65] looks promising to generate high-quality masks in a single step.
Finally, we note that lowering the loss for small timesteps (e.g., $j<25\%$) is not crucial, but speeds-up training by 0.3 to 0.5% PQ. We aim to remove this in future work.
Method # Iters Time [s] PQ [%]
Mask2Former 1 0.02 59.0
LDMSeg 10 0.24 50.0
LDMSeg 20 0.47 50.7
LDMSeg 50 1.14 50.8
Table 1: Inference time to generate an image-conditioned mask on a 4090 GPU. The table provides the results for a different amount of denoising steps. We compare to Mask2Former.
D Limitations and Future Work
Undoubtedly, our model has several limitations despite its general design. We discuss two limitations: (i) the model can miss small background objects due to the projection to latent space; (ii) the model is slower during inference than specialized segmentation models due to the adoption of a diffusion prior. In exchange, our method is simple, general, and unlocks out-of-the-box mask inpainting. Finally, increasing the resolution of the latents, enabling open-vocabulary [55] detection, and including additional dense prediction tasks are exciting directions to explore further.
E Broader Impact
The presented approach relies on pretrained weights from Stable Diffusion. Consequently, our model is subject to the same dataset and architectural biases. The user should be aware of these biases and their impact on the generated masks. For instance, these types of (foundation) models have the ability to hallucinate content.